40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 15 12:28:03 2016, gautam, Update, General, Something has gone wrong - was there a power outage? Vac_15Feb2016.png
    Reply  Mon Feb 15 13:09:33 2016, Koji, Update, General, Something has gone wrong - was there a power outage? 
       Reply  Tue Feb 16 09:13:39 2016, Steve, Update, General, there was a power outage & EQ 20160215_power_outage.png
          Reply  Tue Feb 16 15:02:19 2016, ericq, Update, General, Power Glitch recovery 
             Reply  Tue Feb 16 23:42:22 2016, gautam, Update, General, Power Glitch recovery - arms recovered 
                Reply  Wed Feb 17 13:45:15 2016, ericq, Update, General, No monit on frontends 
             Reply  Thu Feb 25 11:02:36 2016, Steve, Update, General, Power Glitch again powerGlitch.png
                Reply  Thu Feb 25 11:32:04 2016, gautam, Update, General, Power Glitch again 
          Reply  Mon Feb 22 08:45:46 2016, Steve, Update, SUS, EQ 4.3 grabs ITMX-UL magnet EQ4.3LucerneV.pngITMX-UL.png
             Reply  Mon Feb 22 23:00:21 2016, ericq, Update, SUS, ITMX free, green is flashing 
                Reply  Tue Feb 23 09:51:34 2016, Steve, Update, SUS, ITMX free ITMX_FREED.pngITMX_FREE.png
                   Reply  Tue Feb 23 17:46:09 2016, gautam, Update, SUS, EQ  SUS_Summary_23Feb.png
             Reply  Tue Apr 5 08:55:51 2016, Steve, Update, SUS, local EQ 3.1m eq3.1Valencia.png
                Reply  Wed May 25 08:32:55 2016, Steve, Update, SUS, local EQ 3.5m 3.5Cucam.pnglocal3.5cucam.png
                   Reply  Fri Jun 3 14:24:13 2016, Steve, Update, SUS, local EQ 3.1 m eq3.1mMuscoyCa.png
                      Reply  Fri Jun 10 09:52:51 2016, Steve, Update, SUS, EQ 5.2 mag Borego Spring 5.2mBoregoSpring.pngeqBSs.png
                         Reply  Mon Jun 20 09:36:10 2016, Steve, Update, SUS, EQ 3.1 m Big Bear Lake, CA eq3.1BigBearLakeCA.png
Message ID: 12011     Entry time: Thu Feb 25 11:32:04 2016     In reply to: 12010
Author: gautam 
Type: Update 
Category: General 
Subject: Power Glitch again 

 

Quote:

10:15 power glitch today. ETMX Lightwave and air conditions turned back on

The CDS situation was not as catastrophic as the last time, it was sufficient for me to ssh into all the frontends and restart all the models. I also checked that monit was running on all the FEs and that there was no date/time issues like we saw last week. Everything looks to be back to normal now, except that the ntpd process being monitored on c1iscex says "execution failed". I tried restarting the process a couple of times, but each time it returns the same status after a few minutes.

ELOG V3.1.3-